OpenStudy (anonymous):

Log help? What is the solution of log2x-6 256=4?

OpenStudy (anonymous):

X=-5 x=-4 x=4 x=5

OpenStudy (solomonzelman):

\[\log_{2x-6}256=4 ~~~~~~?\]

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

\[\log_AB=C~~~~~~~~~--->~~~~~~~~~~~A^C=B\]

OpenStudy (anonymous):

Is that the change of A rule? Where it makes it, 2x-6^4=256?

OpenStudy (solomonzelman):

\[\log_{2x-6}256=4~~~~~->~~~~(2x-6)^4=256\]

OpenStudy (solomonzelman):

\[(2x-6)^4=256\]\[(2x-6)^4=(~±~4)^2\]\[2x-6=~±~4\] \[2x-6=4~~~~~~~~~~~or~~~~~~~~~~2x-6=-4\]\[2x=10~~~~~~~~~~~~~~~~or~~~~~~~~~~2x=2\]\[x=5~~~~~~~~~~~~~~~~~~~~or~~~~~~~~~~~x=1\]

OpenStudy (anonymous):

Why do you replace the 256?